OpenStudy (anonymous):

the value of y varies directly with x and the constant of variation is 7. what is the value of x when y=63???

jimthompson5910 (jim_thompson5910):

y varies directly with x, so y = k*x for some fixed number k

jimthompson5910 (jim_thompson5910):

we are told that the constant of variation is 7, so k = 7

jimthompson5910 (jim_thompson5910):

we want to find the value of x when y = 63, so, y = k*x y = 7*x 63 = 7*x solve for x